Show how 6 children can share 10 pizzas among themselves

Your illustration should lead you to the following answer 1 4/6


1
Expert's answer
2021-07-09T03:23:36-0400

The number of pizzas are 10 to be shared among 6 children.


When each child gets a pizza, 4 will remain to be shared among them all (6).


Therefore, each child should get 1 and "\frac{4}{6}" .


In other words; "\frac{10}{6}"


=1"\frac{4}{6}" .